OpenStudy (anonymous):

solving systems of Equations by Elimination method: Tickets for a movie cost $5 for adults and $2 for students. One afternoon 21 tickets were sold and the receipts totled $72. How many of each type of ticket was sold? a) write a system of equations. b) solve by using the elimination method. HELP!!!!!!

OpenStudy (saifoo.khan):

Wanna learn how to do this?

hero (hero):

x + y = 21 5x + 2y = 72

OpenStudy (anonymous):

yes!!!!

OpenStudy (saifoo.khan):

Let's suppose the: Adult tickets = x Child ticket = y To find the total tickets sold we will add them. (same as when u have to find the total of something, u add them) So here too we will add them, x + y = Total tickets, therefore, x + y = 21.

OpenStudy (anonymous):

okay got that so far.

OpenStudy (saifoo.khan):

Now for the second equation, (let's say one pastry costs about $5. And u want 8 of them, so it means you will write as, (#of pastries * cost of one pastry) 5(8) for the total amount). So here we have total amount for tickets sold. We will write as, (# of adult tickets *cost of one ticket) + (#of child tickets * cost of one ticket) = Total cost.

OpenStudy (saifoo.khan):

So we have, 5x + 2y = 72

OpenStudy (saifoo.khan):

Finally two equations, x + y = 21 5x + 2y = 72

OpenStudy (anonymous):

okay i got the system of equation, but how do you solve it by using the elimination method can you tech me that?

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=hJOqGkWElNs

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=Y6JsEja15Vk

OpenStudy (saifoo.khan):

I am sure those will help!

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

Welcome.